find equivalent fractions: 6/10, 3/4, 4/12 and 1/5
Answer #1
Multiply or divide the numerator and denominator by the same number.

6/10 = 3/5 = 9/15 = 30/50 and so on.

Do the same with the other fractions.
answered by: R_t